# Bounds on the number of squares in recurrence sequences

### v1

@@ 4.5. $r_{0}>1$ and $\zeta_{4}p/q=p_{r_{0}}/q_{r_{0}}$ for some $4$-th root of unity, $\zeta_{4}$
$1.73r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}$
@@
$1.73r_{0}^{1/2}>\left(\frac{11.46}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}^{2}f_{\ell}^{2}}}\right)^{4}c^{4}\left(\frac{0.1832^{4}|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{4\left|N_{\alpha}\right|^{3}}\right)^{r_{0}}d^{r_{0}}y_{k}^{2r_{0}-2}.$
We can show that $0.1832^{4r_{0}}/r_{0}^{1/2}>0.175^{4r_{0}}$, with the minimum being attained at $r_{0}=3$. Applying this, along with $c=0.75$, yields
$1>\frac{0.739d}{f_{k}^{2}f_{\ell}^{2}\left|N_{\alpha}\right|}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
@@
(4.22) $1>\frac{0.739}{f_{k}^{2}f_{\ell}^{2}}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
@@
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b). Also, as $r_{0}-1\geq 1$, it follows that $0.739\cdot 0.0002344^{r_{0}-1}>\left(0.0001732\right)^{r_{0}-1}$, the above inequality implies
$1>\left(\frac{0.0001732|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}>\left(y_{k}^{2}\frac{d\left(|g|\mathcal{N}_{d,4}\right)^{4}}{5774\left|N_{\alpha}\right|^{3}}\right)^{r_{0}-1}.$
@@
(4.23) $y_{k}\geq\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d,4}\right)^{2}},$
@@ 4.6. Small $d$
If $y_{k}=y^{2}$ is a square, then from (1.1), we know that $x_{k}^{2}-dy^{4}=N_{\alpha}$. By Theorem 1.1 of [1], our theorem holds for $N_{\alpha}=-1$, so we may also assume that $N_{\alpha}\leq-4$. There are 74 such pairs $(a,d)$.
For each such pair, we solved $x^{2}-dy^{4}=N_{\alpha}$ using Magma (version V2.28-2) [8] and its IntegralQuarticPoints() function. For the 74 equations, this calculation took 17.66 seconds using MAGMA’s online calculator. 21 of the equations had at least two solutions in positive integers. Only two of these had three solutions in positive integers:
@@
For $x^{2}-17y^{4}=-16$, note that $\left(103+5^{2}\sqrt{17}\right)/\left(1+\sqrt{17}\right)=\left(161+39\sqrt{17}\right)/8$, so the solutions $(103,5)$ and $(1,1)$ arise from different sequences. Hence Proposition 4.1 holds for $(a,d)=(1,17)$. Similarly, Proposition 4.1 holds for $(a,d)=(2,68)$ too. This completes the proof of Proposition 4.1.
@@
If $N_{\alpha}$ is even, then $|g|\mathcal{N}_{d^{\prime},4}\geq 4$ by Lemma 3.11. So the right-hand side of the inequality in Proposition 4.1 is at most $4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. But for $u\geq 5$, we have $y_{k}>25\left|N_{\alpha}\right|/4$, by Lemma 3.5(c). So Theorem 1.4 holds for $u\geq 5$ and we need only consider $1\leq u\leq 4$.
Similarly, if $N_{\alpha}$ is odd and $u\geq 9$, then we have $y_{k}>81\left|N_{\alpha}\right|/4>19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$ and we need only consider $1\leq u\leq 8$.
@@
Suppose that $y_{k}$ is a square for some $|k|\geq 2$.
From Lemma 3.5(c), $\left|N_{\alpha}\right|=d-a^{2}<d$ and that by Theorem 1.1 of [1] we may assume $\left|N_{\alpha}\right|\geq 4$, we have $y_{k}>\left|N_{\alpha}\right|^{2}u^{4}/8$. So, for $u\geq 2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}\geq 8\left|N_{\alpha}\right|>8\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
However, we saw above that for $N_{\alpha}$ even, if $y_{k}>4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$, there can be no square $y_{\ell}$, satisfying $y_{\ell}>y_{k}>1$. So we need only consider $u=1$ when $N_{\alpha}$ is even.
Similarly, for $N_{\alpha}$ odd, if $u\geq 2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}\geq 18\left|N_{\alpha}\right|>18\left|N_{\alpha}\right|^{2}>90\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. But, we saw above that for $N_{\alpha}$ even, if $y_{k}>19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$, there can be no square $y_{\ell}$, satisfying $y_{\ell}>y_{k}>1$. So we need only consider $u\leq 2$ when $N_{\alpha}$ is odd.
For $u=2$, we have $y_{k}>2\left|N_{\alpha}\right|^{2}$ for $|k|\geq 2$. So if $\left|N_{\alpha}\right|\geq 11$ is odd, then we have $y_{k}\geq 22\left|N_{\alpha}\right|>20\left|N_{\alpha}\right|$. So, for odd $N_{\alpha}$, we need only consider $N_{\alpha}=-9$. In this case, $N(\varepsilon)=\left(t^{2}-4d\right)/4=\pm 1$, so $t^{2}-4a^{2}=-4N_{\alpha}\pm 4=32,40$. This means that $\left(t,a,N_{\alpha}\right)=(\pm 6,1,-9)$. Here we have $y_{\pm 2}\geq 493>20\left|N_{\alpha}\right|$.
For $u=1$, we proceed similarly. Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/8$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $N_{\alpha}$ is even and $9\leq\left|N_{\alpha}\right|\leq 121$ when $-N_{\alpha}$ is an odd square.
For each value of $N_{\alpha}$, the norm of $\varepsilon$ leads us to an equation of the form $t^{2}-a^{2}=-N_{\alpha}\pm 4$. We solve each of these equations over the integers using PARI/GP and calculate $y_{k}$ and $y_{-k}$ for $k\geq 2$ until the lower bound from Proposition 4.1 is exceeded (we never had to go beyond $k=7$). No squares were found. ∎
@@
Suppose that $N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
@@
If $-N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $\gcd\left(a^{2},d\right)=1$ and $u=2$, $t^{2}-du^{2}=-4$.
@@
(1) We start by considering $t^{2}-du^{2}=4$.
@@
(1-i) First, we eliminate $u=4$. We can write $N_{\alpha}=-(2n+1)^{2}$, so $t^{2}-\left(a^{2}+(2n+1)^{2}\right)u^{2}=4$ becomes $t^{2}-64n^{2}-64n=16a^{2}+20$. But $t^{2}\equiv 20\pmod{64}$, has no solution, so $a$ must be odd. Hence $t^{2}-64n^{2}-64n=16a^{2}+20$ implies $t^{2}\equiv 36\pmod{64}$. So $t\equiv\pm 6\pmod{16}$.
@@
(1-ii) Now we eliminate $u=8$. Arguing modulo $9$, from $t^{2}-64d=4$, we see that $d\equiv 0,3,5,6\pmod{9}$. From this, $-N_{\alpha}=d-a^{2}$ being a square and the squares modulo $9$ being $0,1,4,7$, we must have $\left(a^{2}\bmod 9,d\bmod 9\right)=(0,0)$, $(1,5)$, $(4,5)$ or $(7,5)$.
We will show that $d\equiv 5\pmod{9}$ is not possible. In this case, from $t^{2}-64d=4$, we find that $t^{2}\equiv 0\pmod{9}$. So $t\equiv 0\pmod{3}$. Since $d\equiv 5\pmod{9}$), we have $y_{\pm 1}=1+32d\pm 4at\equiv 2\pm 4at\pmod{3}$. Since $3|t$, it follows that $y_{\pm 1}\equiv 2\pmod{3}$, which can never be square. Hence $9|\gcd\left(a^{2},d\right)$.
@@
Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 36$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}<(76/36)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}>16\left|N_{\alpha}\right|$ here. So the case where $u=8$, $-N_{\alpha}$ is an odd square and $t^{2}-du^{2}=4$ is excluded.
@@
(2-ii) Finally, we show that $\gcd\left(a^{2},d\right)=1$ in this case.
We suppose otherwise. since $N_{\alpha}$ is odd, $\gcd\left(a^{2},d\right)$ must be odd and since $d=\left(t^{2}+4\right)/4$, any odd prime factor, $p$, of the gcd must also be a factor of $t^{2}+4$. This means that $t^{2}\equiv-4\pmod{p}$. Hence $p\equiv 1\pmod{4}$. Since $p^{2}|a^{2}$ and $-N_{\alpha}=d-a^{2}$ is a perfect square, this also implies that $p^{2}|d$. Therefore by Lemma 3.11, we have $\left(|g|\mathcal{N}_{d,4}\right)^{2}\geq 100$. Hence the theorem holds if $y_{\pm 1}>(76/100)\left|N_{\alpha}\right|$, by Proposition 4.1. But by Lemma 3.5(c), we know that $y_{\pm 1}>\left|N_{\alpha}\right|$ here. So the theorem holds if $u=2$, $-N_{\alpha}$ is an odd square and $\gcd\left(a^{2},d\right)>1$. ∎
@@
If $-N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.

### v2

@@ 4.5. $r_{0}>1$ and $\zeta_{4}p/q=p_{r_{0}}/q_{r_{0}}$ for some $4$-th root of unity, $\zeta_{4}$
$1.734r_{0}^{1/2}\left(4\frac{d}{\left|N_{\alpha}\right|}\right)^{r_{0}}y_{k}^{2r_{0}+1}>\left(\frac{11.47}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832|g|\mathcal{N}_{d^{\prime},4}}{\left|N_{\alpha}\right|}\right)^{4r_{0}}d^{2r_{0}}y_{k}^{4r_{0}-1}$
@@
$1.734r_{0}^{1/2}>\left(\frac{11.47}{|g|\mathcal{N}_{d^{\prime},4}}\sqrt{\frac{\left|N_{\alpha}\right|}{f_{k}f_{\ell}}}\right)^{4}c^{4}\left(\frac{0.1832^{4}|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{4\left|N_{\alpha}\right|^{3}}\right)^{r_{0}}d^{r_{0}}y_{k}^{2r_{0}-2}.$
We can show that $0.1832^{4r_{0}}/r_{0}^{1/2}>0.175^{4r_{0}}$, with the minimum being attained at $r_{0}=3$. Applying this, along with $c=0.75$ and collecting the terms taken to the power $r_{0}-1$, yields
$1>\frac{0.7405d}{f_{k}^{2}f_{\ell}^{2}\left|N_{\alpha}\right|}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
@@
(4.22) $1>\frac{0.7405}{f_{k}^{2}f_{\ell}^{2}}\left(\frac{0.0002344|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}.$
@@
Since $-N_{\alpha}$ is a square, we have $f_{k}=f_{\ell}=1$ from Proposition 3.1(b). Also, as $r_{0}-1\geq 1$, it follows that $0.7405\cdot 0.0002344^{r_{0}-1}>0.0001735^{r_{0}-1}$. So the above inequality implies
$1>\left(\frac{0.0001735|g|^{4}\mathcal{N}_{d^{\prime},4}^{4}}{\left|N_{\alpha}\right|^{3}}dy_{k}^{2}\right)^{r_{0}-1}>\left(y_{k}^{2}\frac{d\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{4}}{5764\left|N_{\alpha}\right|^{3}}\right)^{r_{0}-1}.$
@@
(4.23) $y_{k}\geq\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}},$
@@ 4.6. Small $d$
If $y_{k}=y^{2}$ is a square, then from (1.1), we know that $x_{k}^{2}-dy^{4}=N_{\alpha}$. By Theorem 1.1 of [1], our theorem holds for $N_{\alpha}=-1$, so we may also assume that $N_{\alpha}\leq-4$. There are 59 such pairs $(a,d)$.
For each such pair, we solved $x^{2}-dy^{4}=N_{\alpha}$ using Magma (version V2.28-2) [4] and its IntegralQuarticPoints() function. For the 59 equations, this calculation took 17.66 seconds using MAGMA’s online calculator. 21 of the equations had at least two solutions in positive integers. Only two of these had three solutions in positive integers:
@@
For $x^{2}-17y^{4}=-16$, since $\left(103+5^{2}\sqrt{17}\right)/\left(1+\sqrt{17}\right)=\left(161+39\sqrt{17}\right)/8$, the solutions $(103,5)$ and $(1,1)$ arise from different sequences. Hence Proposition 4.1 holds for $(a,d)=(1,17)$. Similarly, Proposition 4.1 holds for $(a,d)=(2,68)$ too. This completes the proof of Proposition 4.1.
@@
If $N_{\alpha}$ is even, then $|g|\mathcal{N}_{d^{\prime},4}\geq 4$ by Lemma 3.12. So the right-hand side of the inequality in Proposition 4.1 is $\max\left(1,4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$. But for $u\geq 5$ and $k\neq 0$, we have $y_{k}\geq 25\left|N_{\alpha}\right|/4$ by Lemma 3.5(c). So Theorem 1.4 holds for $u\geq 5$ and we need only consider $1\leq u\leq 4$.
Similarly, if $N_{\alpha}$ is odd and $u\geq 9$, then we have $y_{k}\geq 81\left|N_{\alpha}\right|/4>\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$ for $k\neq 0$ and we need only consider $1\leq u\leq 8$.
Next we treat the case when $K<-1$, where, as in Lemma 3.5, $K$ is the largest negative integer such that $y_{K}>1$.
Lemma 5.1.
If $K<-1$, then there are at most two distinct integer squares among the $y_{k}$’s.
Proof.
From Lemma 3.6, we know that $K<-1$ can only happen in the following two cases:
(i) $a\geq 1$, $d=a^{2}+4$, $t=a$ and $u=1$, where $\alpha=2\varepsilon$ and $N_{\alpha}=-4$,
(ii) $a\geq 1$, $d=a^{2}+1$, $t=2a$ and $u=2$, where $\alpha=\varepsilon$ and $N_{\alpha}=-1$.
In case (i), we have $x_{k}^{2}-dy_{k}^{2}=-4$.
If $d=a^{2}+4$ is even, then it is divisible by $4$ and $x_{k}$ is also even. So the equation becomes $\left(x_{k}/2\right)^{2}-(d/4)y_{k}^{2}=-1$. From Theorem 1.1 of [1], there are at most two distinct squares among the $y_{k}$’s here, so the lemma holds in this case.
If $d$ is odd, the conditions in Theorem 1 of [9] hold with $A=d=a^{2}+4$, $B=1$ both odd and the minimal solution of equation (1) there in odd positive integers being $(1,a)$. Hence there are at most two distinct squares among the $y_{k}$’s in this case too.
In case (ii), we can also apply Theorem 1.1 of [1]. ∎
For the remainder of this section, we may assume that $K=-1$.
@@
From Theorem 1.1 of [1], we cannot have three distinct squares among the $y_{k}$’s for $N_{\alpha}=-1$, so we may assume $\left|N_{\alpha}\right|\geq 4$ for $N_{\alpha}$ even and $\left|N_{\alpha}\right|\geq 9$ for $N_{\alpha}$ odd.
We will suppose that $y_{k}$ is a square with $|k|>1$, but show that this is not possible.
From Lemma 3.5(c) and $\left|N_{\alpha}\right|=d-a^{2}<d$, we obtain $y_{k}\geq\left|N_{\alpha}\right|^{2}u^{4}/10$. So, for $u\geq 2$ and $\left|N_{\alpha}\right|\geq 4$, we have $y_{k}\geq(8/5)\left|N_{\alpha}\right|^{2}\geq(32/5)\left|N_{\alpha}\right|>(32/5)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
However, we saw at the start of this section that for $N_{\alpha}$ even, there can be no squares $y_{k}$ and $y_{\ell}$ satisfying $y_{\ell}>y_{k}>\max\left(1,4.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$. So we need only consider $u=1$ when $N_{\alpha}$ is even.
Similarly, if $-N_{\alpha}$ is an odd square with $N_{\alpha}\leq-9$ and $u\geq 3$, then we have $y_{k}>(81/10)\left|N_{\alpha}\right|^{2}\geq(729/10)\left|N_{\alpha}\right|>(729/10)\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. But, we saw at the start of this section that for $N_{\alpha}$ odd, there can be no squares $y_{k}$ and $y_{\ell}$ satisfying $y_{\ell}>y_{k}>\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$. So we need only consider $u\leq 2$ when $N_{\alpha}$ is odd.
For $u=2$, we have $y_{k}>(8/5)\left|N_{\alpha}\right|^{2}$ for $|k|\geq 2$. So if $\left|N_{\alpha}\right|\geq 25$ is odd, then we have $y_{k}\geq(200/5)\left|N_{\alpha}\right|>40\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$. So, for odd $N_{\alpha}$, we need only consider $N_{\alpha}=-9$. In this case, $N_{\varepsilon}=\left(t^{2}-4d\right)/4=\pm 1$, so $t^{2}-4a^{2}=-4N_{\alpha}\pm 4=32,40$. This means that $\left(t,a,d\right)=(6,1,10)$ (recall from Subsection 1.2 that we only consider positive values of $t$ and $u$). Here we have $y_{\pm 2}\geq y_{-2}=493>19.75\left|N_{\alpha}\right|^{3/2}/\sqrt{d}$.
For $u=1$, we proceed similarly. Here we have $y_{k}>\left|N_{\alpha}\right|^{2}/10$ for $|k|\geq 2$, so we need to consider $4\leq\left|N_{\alpha}\right|\leq 36$ when $-N_{\alpha}$ is an even square and $9\leq\left|N_{\alpha}\right|\leq 169$ when $-N_{\alpha}$ is an odd square.
For each value of $N_{\alpha}$, the norm of $\varepsilon$ leads us to an equation of the form $t^{2}-a^{2}=-N_{\alpha}\pm 4$. We solve each of these equations over the integers using PARI/GP and calculate $y_{k}$ and $y_{-k}$ for $k\geq 2$ until the lower bound from Proposition 4.1 is exceeded (we never had to go beyond $|k|=5$). No squares were found. ∎
@@
If both $y_{\pm 1}$ are squares, then $y_{k}$ is not a square for any $|k|>1$.
Proof.
From Lemmas 3.8(a) and  3.5(c), we have
$y_{1}>\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}y_{-1}\geq\frac{57.32d^{2}}{\left|N_{\alpha}\right|^{2}}\left(\frac{\left|N_{\alpha}\right|u^{2}}{4}\right)^{3}=\frac{57.32d^{2}\left|N_{\alpha}\right|u^{6}}{64}.$
By Proposition 4.1 and Lemma 3.12, if the quantity on the right-hand side exceeds $\max\left(1,19\left|N_{\alpha}\right|^{3/2}/\sqrt{d}\right)$, then the lemma follows. Since $d>\left|N_{\alpha}\right|$, this holds if $d^{2}u^{6}>19\cdot 64/57.32$, which holds unless $d^{2}<21.3$ and $u=1$. This leaves $d=2$ or $3$, but for neither of these does $u=1$ yield a unit. ∎
Lemma 5.4.
Suppose that $-N_{\alpha}$ is an odd square and $y_{\pm 1}$ is a square integer.
@@
If $-N_{\alpha}$ is an odd square, $1\leq u\leq 8$ and $y_{\pm 1}\in\mathbb{Z}$ a square with $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=2$, $t^{2}-du^{2}=-4$ and $\gcd\left(a^{2},d\right)=1$.
@@
We proceed in steps.
(1) We start by showing that $t^{2}-du^{2}=4$ is not possible.
@@
(1-i) First, we eliminate $u=4$.
We can write $N_{\alpha}=-(2n+1)^{2}$, so $t^{2}-\left(a^{2}+(2n+1)^{2}\right)u^{2}=4$ becomes $t^{2}-64n^{2}-64n=16a^{2}+20$. But $t^{2}\equiv 20\pmod{64}$, has no solution, so $a$ must be odd. Hence $t^{2}-64n^{2}-64n=16a^{2}+20$ implies $t^{2}\equiv 36\pmod{64}$. So $t\equiv\pm 6\pmod{16}$.
@@
(1-ii) Now we eliminate $u=8$.
Arguing modulo $9$, from $t^{2}-64d=4$, we see that $d\equiv 0,3,5,6\pmod{9}$. From this, $-N_{\alpha}=d-a^{2}$ being a square and the squares modulo $9$ being $0,1,4,7$, we must have $\left(a^{2}\bmod 9,d\bmod 9\right)=(0,0)$, $(1,5)$, $(4,5)$ or $(7,5)$.
We will show that $d\equiv 5\pmod{9}$ is not possible. In this case, from $t^{2}-64d=4$, we find that $t^{2}\equiv 0\pmod{9}$. So $t\equiv 0\pmod{3}$. Since $d\equiv 5\pmod{9}$, we have $y_{\pm 1}=1+32d\pm 4at\equiv 2\pm 4at\pmod{3}$. Since $3|t$, it follows that $y_{\pm 1}\equiv 2\pmod{3}$, which can never be square. Hence $9|\gcd\left(a^{2},d\right)$.
@@
Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 36$. Hence $\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}<(76/36)\left|N_{\alpha}\right|$. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq 16\left|N_{\alpha}\right|$ here. So the case where $u=8$, $-N_{\alpha}$ is an odd square and $t^{2}-du^{2}=4$ is excluded.
@@
(2-ii) We show that $\gcd\left(a^{2},d\right)=1$ when $u=2$.
We suppose otherwise. Since $N_{\alpha}$ is odd, $\gcd\left(a^{2},d\right)$ must be odd and since $d=\left(t^{2}+4\right)/4$, any odd prime factor, $p$, of the gcd must also be a factor of $t^{2}+4$. This means that $t^{2}\equiv-4\pmod{p}$. Hence $p\equiv 1\pmod{4}$. Since $p^{2}|a^{2}$ and $-N_{\alpha}=d-a^{2}$ is a perfect square, this also implies that $p^{2}|d$. Therefore by Lemma 3.12, we have $\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}\geq 100$. Hence the theorem holds if $y_{\pm 1}>(76/100)\left|N_{\alpha}\right|$, by Proposition 4.1. But by Lemma 3.5(c), we know that $y_{\pm 1}\geq\left|N_{\alpha}\right|$ here. So the theorem holds if $u=2$, $-N_{\alpha}$ is an odd square and $\gcd\left(a^{2},d\right)>1$. ∎
@@
If $-N_{\alpha}$ is an even square, $1\leq u\leq 4$ and $y_{\pm 1}<\max\left(1,\frac{76\left|N_{\alpha}\right|^{3/2}}{\sqrt{d}\left(|g|\mathcal{N}_{d^{\prime},4}\right)^{2}}\right)$, then we must have $u=1$, $t^{2}-du^{2}=-4$, $N_{\alpha}\equiv 12\pmod{16}$ and $\gcd\left(a^{2},d\right)=1,4$.